OpenStudy (anonymous):

how do i solve this? x^2 -4/x^2-1 multiplied by x+1/x+2

OpenStudy (helder_edwin):

do you have to simplify?

OpenStudy (anonymous):

yes, i have to multiply and simplify

OpenStudy (zeesbrat3):

the answer is 2 if I did my math right

OpenStudy (helder_edwin):

No you didn't @zeesbrat3

OpenStudy (helder_edwin):

actually it would be better if you factor the first fraction yo have do you know how to do this?

OpenStudy (zeesbrat3):

Oh well, at least I tried. Then again maybe I shouldve written it down...

OpenStudy (anonymous):

\[\huge \frac{x^2-4}{x^2-1} \times \frac{x+1}{x+2}\] \[\huge \frac{(x+2)(x-2)}{(x+1)(x-1)}\times \frac{x+1}{x+2}\]

OpenStudy (lgbasallote):

^best possible explanation :)

OpenStudy (anonymous):

nphuongsun93 how did u do that?^^^

OpenStudy (anonymous):

\[\huge x^2-4 = (x+2)(x-2)\] \[\huge x^2-1 = (x+1)(x-1)\]

OpenStudy (anonymous):

formula \[\huge a^2 - b^2 = (a-b)(a+b)\]

OpenStudy (anonymous):

can u help me with my questions plz?

OpenStudy (helder_edwin):

@nique504 did you understand what @nphuongsun93 did? what is the answer?

OpenStudy (anonymous):

i think so, the answer would be x-2/x-2 ?

OpenStudy (anonymous):

opps i meant x-2/x-1

OpenStudy (helder_edwin):

YES!

OpenStudy (anonymous):

yayy! haha thank you all for helping :)